Grade 10Electromagnetic Induction

How many time constants will elapse the current in a charging RC circuit drops to half of its initial value? Answer the same question for a discharging RC circuit.

Hrishant Goswami
12 Years agoGrade 10

1 Answer

Jitender Pal
12 Years ago
Sol. Q/2 = Q(1 – e^-t/CR) ⇒ 1/2 = (1 – e^-t/CR) ⇒ e^-t/CR = ½ ⇒ t/RC = log2 ⇒ n = 0.69.